OpenStudy (kaliyah0524):

Will fan & medal ! 1. (3x + 1/2) (3x - 1/2) 2. (12x^3) (1/2x^6) 3.[2+(4b+4)]^2 4.[(9s+4)+8][(9s+4)-8]

OpenStudy (danjs):

what do you need to do, just expand those out and simplify it?

OpenStudy (kaliyah0524):

For 1 & 2 you FOIL it. I know how to do it, i just dont work with fractions that good

OpenStudy (kaliyah0524):

For 3 & 4, i dont know because i was absent when it was taught @DanJS

OpenStudy (danjs):

Yeah, need to distribute the first one, multiply like this pattern (a + b)*(c + d) = ac + ad + bc + bd

OpenStudy (kaliyah0524):

Yeah i know how to foil.. Its the fraction part i cant do

OpenStudy (legomyego180):

\[(\frac{ 1 }{ 2 })(\frac{ 3 }{ 1 })=\frac{ 6 }{ 2 }=3\]

OpenStudy (legomyego180):

It's okay, fractions are hard. Even in some of the higher level maths one of the most common mistakes made is dealing with these same rules with fractions.

OpenStudy (kaliyah0524):

Thanks

OpenStudy (danjs):

yeah multiplying fractions is easier than adding since you dont need a common denominator \[\large \frac{ a }{ b }*\frac{ c }{ d } = \frac{ a*c }{ b*d }\]

OpenStudy (legomyego180):

Right, exactly.

OpenStudy (kaliyah0524):

Got it. Thank you guys